# MUSE-Net: Missingness-aware mUlti-branching Self-attention Encoder for Irregular Longitudinal Electronic Health Records

### v1

@@ 8.3 Algorithm for GPU Acceleration in MGP Imputation
$=$ $\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}\left[\text{Tr}\left((\Sigma_{p}^{o})^{-1}\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\mathcal{P}^{-1}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}\right)\right]=\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}\left[\boldsymbol{d}_{i}^{T}(\Sigma_{p}^{o})^{-1}\left(\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\mathcal{P}^{-1}\boldsymbol{d}_{i}\right)\right]$
@@
where $\boldsymbol{d}_{i}^{T}(\Sigma_{p}^{o})^{-1}\approx([U_{k}]_{:,i+1})^{T}$ are computed by mPCG, and $\mathcal{P}^{-1}$ can be calculated using the Matrix Inversion Lemma, i.e., $\mathcal{P}^{-1}=(C_{0}C_{0}^{T}+E)^{-1}=E^{-1}-E^{-1}(I+C_{0}^{T}E^{-1}C_{0})^{-1}C_{0}^{T}E^{-1}$, which avoids the inversion computation of the $O_{p}\times O_{p}$ matrix using the inversion of a smaller $r_{0}\times r_{0}$ matrix. To estimate $\log|\Sigma_{p}^{o}|$, we integrate stochastic trace estimation with partial Lanczos tridiagonalizations [59, 61], i.e., $\mathcal{T}_{1},\mathcal{T}_{2},\cdots,\mathcal{T}_{t}$, from mPCG. The key idea of partial Lanczos tridiagonalizations is that instead of calculating the full tridiagonalization, i.e., $(C^{-1})^{T}\Sigma_{p}^{o}C^{-1}=H^{T}\mathcal{T}H$ where $H$ is orthonormal, we run $k$ iterations ($k\ll O_{p}$) of this algorithm $t$ times so that we obtain $t$ partial Lanczos tridiagonalizations decompositions $\mathcal{T}_{i}$ (see line 13 and 14 in Algorithm 1). Specifically, the log determinant $\log|\Sigma_{p}^{o}|$ can be calculated by:
$\log|\Sigma_{p}^{o}|$ $=$ $\log|\Sigma_{p}^{o}C^{-1}\mathcal{P}(C^{-1})^{T}|=\log|(C^{-1})^{T}\Sigma_{p}^{o}C^{-1}|+\log|\mathcal{P}|$ (27)
$\log|(C^{-1})^{T}\Sigma_{p}^{o}C^{-1}|$ $=$ $\text{Tr}(\log\mathcal{T})=\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}[(\boldsymbol{d}_{i}^{T}(C^{-1})^{T})(\log\mathcal{T})(C^{-1}\boldsymbol{d}_{i})]$ (28)
$=$ $\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}[(\boldsymbol{d}_{i}^{T}(C^{-1})^{T})H_{i}^{T}(\log\mathcal{T}_{i})H_{i}(C^{-1}\boldsymbol{d}_{i})]\approx\frac{1}{t}\sum_{i=1}^{t}\boldsymbol{e}_{1}^{T}(\log T_{i})\boldsymbol{e}_{1}$
@@
The calculated $(\Sigma_{p}^{o})^{-1}\boldsymbol{y}_{p}^{o}$, $\log|\Sigma_{p}^{o}|$, and $\text{Tr}\left((\Sigma_{p}^{o})^{-1}\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\right)$ are used for computing $\mathcal{L}(\boldsymbol{\Theta})$ and its derivative $d\mathcal{L}/d\boldsymbol{\Theta}$ as shown in lines 17 and 18. These values will be further utilized to estimate the optimal MGP hyperparameters through a gradient-based optimization algorithm (i.e., ADAM Optimizer [63]), denoted as $\boldsymbol{\Theta}^{*}$, which is subsequently utilized to compute the posterior mean, completing data imputation as shown in line 22.

### v2

@@ VI-C Algorithm for GPU Acceleration in MGP Imputation
$=$ $\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}\left[\text{Tr}\left((\Sigma_{p}^{o})^{-1}\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\mathcal{P}^{-1}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}\right)\right]$
$=$ $\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}\left[\boldsymbol{d}_{i}^{T}(\Sigma_{p}^{o})^{-1}\left(\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\mathcal{P}^{-1}\boldsymbol{d}_{i}\right)\right]$
@@
where $\boldsymbol{d}_{i}^{T}(\Sigma_{p}^{o})^{-1}\approx([U_{k}]_{:,i+1})^{T}$ are computed by mPCG, and $\mathcal{P}^{-1}$ can be calculated using the Matrix Inversion Lemma, i.e., $\mathcal{P}^{-1}=(C_{0}C_{0}^{T}+E)^{-1}=E^{-1}-E^{-1}(I+C_{0}^{T}E^{-1}C_{0})^{-1}C_{0}^{T}E^{-1}$, which avoids the inversion computation of the $O_{p}\times O_{p}$ matrix using the inversion of a smaller $r_{0}\times r_{0}$ matrix. To estimate $\log|\Sigma_{p}^{o}|$, we integrate stochastic trace estimation with partial Lanczos tridiagonalizations [60, 62], i.e., $\mathcal{T}_{1},\mathcal{T}_{2},\cdots,\mathcal{T}_{t}$, from mPCG. The key idea is that instead of calculating the full tridiagonalization, i.e., $(C^{-1})^{T}\Sigma_{p}^{o}C^{-1}=H^{T}\mathcal{T}H$ where $H$ is orthonormal, we run $k$ iterations ($k\ll O_{p}$) of this algorithm $t$ times so that we obtain $t$ partial Lanczos tridiagonalizations decompositions $\mathcal{T}_{i}$ (see line 13 and 14 in Algorithm 1). Specifically, the log determinant $\log|\Sigma_{p}^{o}|$ can be calculated by:
$\log|\Sigma_{p}^{o}|=\log|\Sigma_{p}^{o}C^{-1}\mathcal{P}(C^{-1})^{T}|$
$=$ $\log|(C^{-1})^{T}\Sigma_{p}^{o}C^{-1}|+\log|\mathcal{P}|$
$\log|(C^{-1})^{T}\Sigma_{p}^{o}C^{-1}|=\text{Tr}(\log\mathcal{T})$
$=$ $\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}[(\boldsymbol{d}_{i}^{T}(C^{-1})^{T})(\log\mathcal{T})(C^{-1}\boldsymbol{d}_{i})]$
$=$ $\underset{\boldsymbol{d}_{i}\sim\mathcal{N}(0,\mathcal{P})}{\mathbb{E}}[(\boldsymbol{d}_{i}^{T}(C^{-1})^{T})H_{i}^{T}(\log\mathcal{T}_{i})H_{i}(C^{-1}\boldsymbol{d}_{i})]$
$\approx$ $\frac{1}{t}\sum_{i=1}^{t}\boldsymbol{e}_{1}^{T}(\log T_{i})\boldsymbol{e}_{1}$ (29)
@@
The calculated $(\Sigma_{p}^{o})^{-1}\boldsymbol{y}_{p}^{o}$, $\log|\Sigma_{p}^{o}|$, and $\text{Tr}\left((\Sigma_{p}^{o})^{-1}\frac{d\Sigma_{p}^{o}}{d\boldsymbol{\Theta}}\right)$ are used for computing $\mathcal{L}(\boldsymbol{\Theta})$ and its derivative $d\mathcal{L}/d\boldsymbol{\Theta}$ (see lines 17 and 18). These values will be further utilized to estimate the MGP hyperparameters through a gradient-based optimization algorithm (i.e., ADAM Optimizer [64]), denoted as $\boldsymbol{\Theta}^{*}$, which is subsequently utilized to compute the posterior mean, completing data imputation as shown in line 22.